# LiLIS: Enhancing Big Spatial Data Processing with Lightweight Distributed Learned Index

### v1

@@ Abstract
The efficient management of big spatial data is crucial for location-based services, particularly in smart cities. However, existing systems such as Simba and Sedona, which incorporate distributed spatial indexing, still incur substantial index construction overheads, therefore rendering them far from optimal for real-time analytics. Recent studies demonstrate that learned indices can achieve high efficiency through a well-designed machine learning model, but how to design learned index for distributed spatial analytics remains unaddressed. In this paper, we present LiLIS, a Lightweight Distributed Learned Index for big Spatial data. LiLIS amalgamates machine-learned search strategies with spatial-aware partitioning within a distributed framework, and efficiently implements common spatial queries, including point query, range query, $k$-nearest neighbors ($k$NN), and spatial joins. Extensive experimental results over real-world and synthetic datasets show that LiLIS outperforms state-of-the-art big spatial data analytics by 2–3 orders of magnitude for most spatial queries, and the index building achieves 1.5-2$\times$ speed-up. The code is available at https://github.com/SWUFE-DB-Group/learned-index-spark.
keywords:
Learned index , Spatial data , Distributed analytics
journal: XXX
[…]
[swufe]organization=Southwestern University of Finance and Economics, city=Chengdu, postcode=611130, country=China
[…]
[cq]organization=Chongqing University, city=Chongqing, postcode=400044, country=China
1 Introduction
Owing to the proliferation of ubiquitous GPS-enabled devices—which not only generate massive spatial data but also enable the collection and sharing of trajectories—from a myriad of location-based services (LBS) such as Google Maps, Uber and geo-tagged social media platforms, researchers now have unprecedented opportunities to analyze movement patterns [1]. Meanwhile, as objects increasingly carry both geographic coordinates and textual descriptors, spatial keyword queries (SKQs)—which jointly optimize spatial proximity and textual relevance—have been extensively studied [2]. For example, DiDi Chuxing [^1] https://web.didiglobal.com/, a leading ride-sharing platform in China, receives more than 25 million trip orders and produces over 15 billion location points per day [3]. To efficiently manage the voluminous spatial datasets, technological innovations in terms of distributed storage and processing have been proposed to handle the limitation of a single-node machine [4, 5]. State-of-the-art big spatial data analytics, such as Simba [6] and Sedona [7], are able to provide high throughput by providing inherent spatial indexing and optimization. Nevertheless, current solutions which rely on traditional spatial indices such R-tree [8] and Quadtree [9] are still hindered by considerable index construction overhead and runtime query cost. Therefore, developing efficient approaches to manage and process large-scale spatial data remains a critical challenge for modern data systems [10].
@@ 1 Introduction
There are several key challenges to design learned spatial indices for distributed analytics. Firstly, training a machine learning model to achieve high-quality predictions is no trivial task, as it often involves a time-consuming training process, especially when dealing with large-scale datasets (C1). Secondly, most model-based learned index methods are not compatible with existing query paradigms, necessitating significant refactoring efforts (C2). Last but not the least, existing learned index methods support only a limited range of spatial queries, restricting their applicability in diverse real-world scenarios (C3). To bridge the gap mentioned above, we present a Lightweight Distributed Learned Index for big Spatial data, dubbed as LiLIS. Specifically, In order to address those challenges, several novel techniques in terms of distributed learned spatial index are developed in this paper:
  • 1.
    Inspired by RadixSpline [18], we leverage an error-bounded spline interpolation to implement LiLIS by learning the two-dimensional distribution of the underlying spatial data. This learned index can be realized with very few parameters using simple yet effective interpolating techniques, making the spatial index considerable lightweight compared to traditional spatial indices (e.g., R-tree and KD-tree), and it also serves as the foundation for efficient spatial queries. (C1)
  • 2.
    Secondly, unlike other learned indices, LiLIS is compatible with existing execution engines of big data systems, and thus provides native features such as scalability and fault tolerance. For example, LiLIS can utilize spatial-aware partitioning and two-phrase pruning to boost the performance. The LiLIS prototype implemented on Apache Spark is straightforward without hacking the underlying execution engine. (C2)
  • 3.
@@
The remainder of this paper is organized as follows: Section 2 reviewed the related work. The overview design and key methodology of LiLIS are described in Section 3. We discussed the detailed efficient distributed spatial algorithms in Section 4. Section 5 presents the results of our experiments, providing insights about the efficiency and scalability of LiLIS, and Section 6 concludes this paper and suggests directions for future research.
2 Related Work
2.1 Spatial Data Indexing
@@
Conventional spatial indexing schemes mentioned above are typically designed as general-purpose structures without optimizations aligned to particular data distribution patterns or domain-specific characteristics. Consequently, when applied to large-scale, high-dimensional datasets, they frequently suffer from reduced query throughput and incur significant storage and I/O overheads [25, 26, 27].
2.2 Learned Indices
@@
A learned index is a technique that employs machine‑learning models to supplant traditional database indexing structures (e.g., B‑trees and hash tables). The idea behind learned index is to learn a function that maps a search key to the storage address of a data object [12, 28, 29, 30]. T. Kraska et al. [11] first introduced the concept of learned indexes by proposing the Recursive Model Index (RMI), which hierarchically models data distributions through a multi-stage neural network that effectively learns a Cumulative Distribution Function (CDF) to predict the rank of a search key [31]. However, while Kraska et al.’s foundational work demonstrates that a learned index can deliver up to a threefold reduction in search time and an order-of-magnitude decrease in memory footprint [11], it remains confined to static, read-only workloads. Subsequently, J. Ding et al. introduced ALEX [32], an updatable learned index that adaptively partitions data to sustain high query and insert performance under dynamic workloads. Beyond the one-dimensional learned indices, later work further extends them into multi-dimensional settings [33]. The Z-order model [14] extends RMI to spatial data by using a space-filling curve to order points and learning the CDF to map each point’s key to its rank, and the recursive spatial model index (RSMI) [34] further develops these principles by employing the rank space–based transformation technique [35].
@@
Although research on learned indices has significantly enhanced query performance, current studies are confined to single-machine environments and fall short in addressing large-scale data processing. At present, there is a significant research gap in the field of distributed learned indices [36].
2.3 Distributed Spatial Analytics
To process large scale spatial data, a large body of big spatial analytics have been proposed over the years [37]. Simba [6], an extension of SparkSQL that incorporates the DataFrame API to enable efficient spatial data queries, features a two-stage indexing design with spatial native query optimizations on resilient distributed datasets (RDDs). SpatialSpark [38] is a framework built on top of Spark RDDs that distributes spatial data across multiple Spark nodes and constructs local R-tree indices on each node, thereby reducing inter-node data transfer, enhancing overall system performance, and supporting both range queries and spatial join queries for large-scale spatial data. GeoMesa [39] is an open-source framework that provides scalable geospatial analytics on distributed infrastructures by implementing spatio-temporal indexing over Apache Accumulo, HBase, Redis, and Cassandra, and integrating near-real-time stream processing via Apache Kafka to support the efficient storage, indexing, and querying of large-scale spatial and spatio-temporal datasets. Apache Sedona [40] is a state-of-the-art cluster computing system that extends engines such as Apache Spark, Apache Flink, and Snowflake with out-of-the-box distributed Spatial Datasets and Spatial SQL to efficiently load, process, and analyze large-scale spatial data across machines.
@@ 2.3 Distributed Spatial Analytics
Existing distributed spatial systems typically rely on traditional indexing methods [17], such as R-trees or grids, which impose considerable I/O and memory overhead during both construction and querying. However, to the best of our knowledge, there is no any distributed spatial analytics equipped with learned spatial indies yet.
3 LiLIS Design
@@
3.1 Spatial-aware Partitioner
@@ 1 $SR\leftarrow$ sampling over $R$;
2 Construct grids list $G$ through the speicified method over $SR$;
@@ 21 Map $(id,o)$ to $o$ over $R^{*}$;
The outer for-loop (Lines 3-15) is indeed a parallel map operation in Apache Spark, and each spatial object $o$ will be mapped into a tuple $(id,o)$, where $id$ is the identifier of the grid in which $o$ locates. The enumerate function borrows from Python [^2] https://docs.python.org/3/library/functions.html#enumerate (Line 5), emitting a tuple containing an $id$ (starting from 0) and the $grid$ obtained from iterating over $G$. As for overflowed object, its partition $id$ will be assigned to the size of grids (Lines 12-14). Therefore, a spatial RDD $R$ will be transformed into a key-value RDD $R^{*}$, and a re-partitioning procedure (Line 16) through shuffling can be conducted over $R^{*}$ by the key (i.e., the $id$ of the partition). Finally, $R^{*}$ will be mapped back into a spatial RDD with objects only (Line 17). Since the partitioner acts as a global index, the master node must maintain all partitions’ properties within $G$, including their spatial boundaries and identifiers.
3.2 Distributed Learned Spatial Index
To keep the compatibility with existing systems [6, 7, 41], LiLIS follows the two-phase filtering solution, and the local index is implemented via a learned model within a given partition. Assume that a spatial object is $(x,y,v)$, where $x$, $y$, and $v$ is its latitude, longitude, and extra attribute, respectively. The two-dimensional spatial points shall be projected into one dimension for sorting, and possible sorting criteria can be either one arbitrary axis or some aggregated value (e.g., Z-order curve and GeoHash). In this paper, we denote the chosen sorting criteria as the key of a spatial object, so each spatial point can be represented as a tuple of $(key,x,y,v)$.
@@ 15       $T[prev+1]\leftarrow n-1$;
Remark 1
It is trivial to convert a string into an unsigned integer because we can apply well-tested hash code computation methods (e.g., product sum) directly as in Java language.
4 Query Algorithms in LiLIS
In this section, we discuss how to design efficient query algorithms based on the learned index proposed in Section 3.
4.1 Point Query
@@ 4.2 Range Query
Remark 2
The range query above is indeed a rectangle range query. As for a circle range query with a center point $p=(x,y)$ and a radius $r$, a practical solution is to construct a minimal bounding rectangle (MBR) for this circle, and perform a regular rectangle range query first. Finally, a subsequent refinement step is necessary to filter out false positives.
4.3 $k$NN Query
@@
$\lceil\frac{\ln{\sqrt{(x_{u}-x_{l})^{2}+(y_{u}-y_{l})^{2}}}-\ln{\sqrt{k(x_{u}-x_{l})(y_{u}-y_{l})/(\pi N)}}}{\ln{\frac{4k}{\pi(k-1)}}}\rceil$ (3)
@@ 4.4 Join Query
Remark 3
LiLIS also support a few other types of joins (e.g., $k$NN-join) out of the box, but how to leverage the learned index to design a unified framework for more types of joins is left as a future work.
5 Experiment
In this section, we conduct extensive experiments to demonstrate the superiority of LiLIS over both real-world and synthetic datasets. All algorithms in this paper were implemented in Java, and all experiments were performed in a cluster with 7 CentOS machines on Tencent Cloud, in which each machine is shipped with 8 GB memory, 80 GB SSD and Intel Duo Processor of 3.0 GHz. As mentioned before, the design of LiLIS is independent with the specific distributed engine, and we present it on Apache Spark 3.0 using Standalone mode, in which one machine is chosen as the master node, and the remaining 6 machines are slave nodes.
5.1 Setup
5.1.1 Datasets
The datasets used in the experiments are summarized in Table 1. Both CHI [^3] https://data.cityofchicago.org/ and NYC [^4] https://www.nyc.gov/site/tlc/about/data.page contain extra information in addition to spatial locations, while SYN consists of spatial points only, which are generated by Spider [^5] https://spider.cs.ucr.edu/. As we can see, although it is possible to load CHI and SYN in a single-node machine, the performance in a distributed setting among different frameworks is a meaningful research question, and the evaluation is enough to demonstrate the effectiveness and efficiency of LiLIS. As for spatial join queries, we constructed 7000 polygons in Chicago, and perform a spatial join with CHI. For other queries, we use NYC as the default dataset.
Table 1: Spatial Datasets
@@ 5.1.1 Datasets
5.1.2 Baselines and Implementations
For evaluation, we compare LiLIS with the state-of-the-art big spatial analytics Sedona [7]. Sedona is shipped with two indices (including R-tree and Quadtree) and two partitioners (including KD-tree and Quadtree), and we use Sedona-{I}{P} to denote its various implementation variants, in which {I} and {P} mean the index type and partitioner, respectively. For example, Sedona-RK uses R-tree as the index while adopts KD-tree as the partitioner. The four variants of Sedona are summarized in Table 2. Notably, Sedona also allows the no-indexing solution, and we donate it as Sedona-N. As for the implementation using vanilla Apache Spark, we donate it as Spark. As discussed in Section 3, LiLIS can adopt different partitioners, we use LiLIS-F, LiLIS-A, LiLIS-Q, LiLIS-K, and LiLIS-R to denote the implementations with fixed-grid, adaptive-grid, quadtree, KD-tree, and R-tree, respectively. In our experiments, we adopt the KD-tree as the default partitioning method for LiLIS. Consequently, LiLIS-K serves as the default implementation in our proposed framework.
Table 2: Sedona Variants

### v3

@@ Abstract
The efficient management of big spatial data is crucial for location-based services, particularly in smart cities. However, existing systems such as Simba and Sedona, which incorporate distributed spatial indexing, still incur substantial index construction overheads, rendering them far from optimal for real-time analytics. Recent studies demonstrate that learned indices can achieve high efficiency through a well-designed machine learning model, but how to design learned index for distributed spatial analytics remains unaddressed. In this article, we present LiLIS, a Lightweight distributed Learned Index for big Spatial data. LiLIS amalgamates machine-learned search strategies with spatial-aware partitioning within a distributed framework, and efficiently implements common spatial queries, including point query, range query, $k$-nearest neighbors ($k$NN), and spatial joins. Extensive experimental results over real-world and synthetic datasets show that LiLIS outperforms state-of-the-art big spatial data analytics by 2-3 orders of magnitude for most spatial queries, and the index building achieves 1.5-2$\times$ speed-up. The code is available at https://github.com/SWUFE-DB-Group/learned-index-spark.
Index Terms:
Learned index, spatial data, spatial databases, distributed analytics.
I Introduction
Tthe widespread adoption of GPS-enabled devices has led to the generation of extensive spatial data, enabling the collection and dissemination of trajectory information through location-based services (LBS) such as Google Maps, Uber, and geo-tagged social media platforms [1, 2, 3, 4]. This development provides researchers with unprecedented opportunities to mine large scale spatial data. For example, DiDi Chuxing [^1] https://web.didiglobal.com/, a leading ride-sharing platform in China, receives more than 25 million trip orders and produces over 15 billion location points per day [5]. To efficiently manage the voluminous spatial datasets, technological innovations in terms of distributed storage and processing have been proposed to handle the limitation of a single-node machine [6, 7]. State-of-the-art big spatial data analytics, such as Simba [8] and Sedona [9], are able to provide high throughput by providing inherent spatial indexing and optimization. Nevertheless, current solutions which rely on traditional spatial indices such R-tree [10] and Quadtree [11] are still hindered by considerable index construction overhead and runtime query cost. Therefore, developing efficient approaches to manage and process large-scale spatial data remains a critical challenge for modern data systems [12].
@@ I Introduction
There are several key challenges to design learned spatial indices for distributed analytics. Firstly, training a machine learning model to achieve high-quality predictions is no trivial task, as it often involves a time-consuming training process, especially when dealing with large-scale datasets (C1). Secondly, most model-based learned index methods are not compatible with existing query paradigms, necessitating significant refactoring efforts (C2). Last but not the least, existing learned index methods support only a limited range of spatial queries, restricting their applicability in diverse real-world scenarios (C3). To bridge the gap mentioned above, we present a Lightweight distributed Learned Index for big Spatial data systems, dubbed as LiLIS. Specifically, In order to address those challenges, several novel techniques in terms of distributed learned spatial index are developed in this paper:
  • We leverage an error-bounded spline interpolation to implement LiLIS by learning the two-dimensional distribution of the underlying spatial data. This learned index can be realized with very few parameters using simple yet effective interpolating techniques, making the spatial index considerable lightweight compared to traditional spatial indices (e.g., R-tree and KD-tree), and it also serves as the foundation for efficient spatial queries. (C1)
  • Unlike other learned indices, LiLIS is compatible with existing execution engines of big data systems, and thus provides native features such as scalability and fault tolerance. For example, LiLIS can utilize spatial-aware partitioning and two-phase pruning to boost the performance. The LiLIS prototype implemented on Apache Spark is straightforward without hacking the underlying execution engine. (C2)
@@
The remainder of this article is organized as follows: Section II reviewed the related work. The overview design and key methodology of LiLIS are described in Section III. We discussed the details of efficient distributed spatial algorithms in Section IV. Section V presents the results of our experiments, providing insights about the efficiency and scalability of LiLIS, and Section VI concludes this article and suggests directions for future research.
II Related Work
II-A Spatial Data Indexing
@@
Conventional spatial indexing schemes mentioned above are typically designed as general-purpose structures without optimizations aligned to particular data distribution patterns or domain-specific characteristics. Consequently, when applied to large-scale, high-dimensional datasets, they frequently suffer from reduced query throughput and incur significant storage and I/O overheads [26, 27, 28]. Therefore, there is a compelling need for a lightweight yet efficient solution to mitigate these challenges.
II-B Learned Indices
@@
A learned index is an advanced methodology that leverages machine learning models to replace conventional database indexing structures, such as B-trees and hash tables. The core principle of a learned index is to develop a function that maps a search key to the corresponding storage address of a data object, thereby optimizing query performance [14, 29, 30, 31]. T. Kraska et al. [13] first introduced the concept of learned indexes, and they proposed the Recursive Model Index (RMI), which hierarchically models data distributions through a multi-stage neural network [32]. However, while their foundational work demonstrates that a learned index can deliver up to a threefold reduction in search time and an order-of-magnitude decrease in memory footprint [13], it remains confined to static, read-only workloads. Subsequently, J. Ding et al. introduced ALEX [33], an updatable learned index that adaptively partitions data to sustain high query and insert performance under dynamic workloads. Beyond the one-dimensional learned indices, later work further extends them into multi-dimensional settings [34]. The Z-order model [16] extends RMI to spatial data. Their method uses a space-filling curve to order points and learns the Cumulative Distribution Function (CDF) to map each point’s key to its rank. The recursive spatial model index [35] further develops these principles by employing the rank space–based transformation technique [36].
@@
Although research on learned indices has significantly enhanced query performance, current studies are confined to single-machine environments and fall short in addressing large-scale data processing. At present, there is a significant research gap in the field of distributed learned indices.
II-C Distributed Spatial Analytics
To process large scale spatial data, a large body of big spatial analytics have been proposed over the years [37, 38]. Simba [8] enhances SparkSQL with DataFrame based spatial queries using two-stage RDD indexing. SpatialSpark [39] implements node-local R-trees on Spark RDDs to minimize data transfer. GeoMesa [40] provides spatio-temporal indexing across multiple database backends with Kafka integration. Among these, Apache Sedona [41] stands as state-of-the-art solution, and it delivers superior performance for large-scale geospatial processing through its innovative distributed indexing and query optimization architecture.
@@ II-C Distributed Spatial Analytics
Existing distributed spatial systems typically rely on traditional indexing methods [19], such as R-trees or grids, which impose considerable I/O and memory overhead during both construction and querying. However, there is no any distributed spatial analytics equipped with learned spatial indies yet.
III LiLIS Design
@@
III-A Spatial-aware Partitioner
@@ 1 $SR\leftarrow$ sampling over $R$;
2 Construct grids list $G$ through the specified method over $SR$;
@@ 21 Map $(id,o)$ to $o$ over $R^{*}$;
The outer for-loop (Lines 3-15) is indeed a parallel map operation in Apache Spark, and each spatial object $o$ will be mapped into a tuple $(id,o)$, where $id$ is the identifier of the grid in which $o$ locates. The enumerate function borrows from Python (Line 5), emitting a tuple containing an $id$ (starting from 0) and the $grid$ obtained from iterating over $G$. As for overflowed object, its partition $id$ will be assigned to the size of grids (Lines 12-14). Therefore, a spatial RDD $R$ will be transformed into a key-value RDD $R^{*}$, and a re-partitioning procedure (Line 16) through shuffling can be conducted over $R^{*}$ by the key (i.e., the $id$ of the partition). Finally, $R^{*}$ will be mapped back into a spatial RDD with objects only (Line 17). Since the partitioner acts as a global index, the master node must maintain all partitions’ properties within $G$, including their spatial boundaries and identifiers.
III-B Distributed Learned Spatial Index
To keep the compatibility with existing systems [8, 9, 42], LiLIS follows the two-phase filtering solution, and the local index is implemented via a learned model within a given partition. Assume that a spatial object is $(x,y,v)$, where $x$, $y$, and $v$ is its latitude, longitude, and extra attribute (e.g., the textural description), respectively. The two-dimensional spatial points shall be projected into one dimension for sorting, and possible sorting criteria can be either one arbitrary axis or some aggregated value (e.g., Z-order curve and GeoHash). In this paper, we denote the chosen sorting criteria as the key of a spatial object, so each spatial point can be represented as a tuple of $(key,x,y,v)$.
@@ 15       $T[prev+1]\leftarrow n-1$;
Remark 1: Converting a string to an unsigned integer is a well-established process, as it can be efficiently accomplished by leveraging robust, thoroughly validated hash code computation techniques, such as the product-sum method, which is natively implemented in programming languages like Java.
III-C Discussions on LiLIS Design
III-C1 Enhancing Rather Than Replacing
Prior work on learned indices often focuses on replacing traditional indexing structures, which necessitates significant refactoring of system architectures and algorithm implementations. In contrast, LiLIS adopts a complementary approach by enhancing the existing computation engine without requiring disruptive overhauls, thus reducing adoption barriers for real-world systems.
III-C2 Lightweight
Motivated by the objective of enhancing efficiency, LiLIS prioritizes a lightweight design. Consequently, constructing a learned index with LiLIS entails minimal overhead due to its reliance on a limited number of parameters. Furthermore, an additional benefit lies in the ability to develop a dual-indexing framework, enabling the selection of an appropriate index based on specific requirements. For instance, when either $x$ or $y$ serves as the key, the construction of an R-tree and a learned index can leverage a shared sorting procedure, thereby optimizing the indexing process.
IV Query Algorithms in LiLIS
In this section, we discuss how to design efficient query algorithms based on the learned index proposed in Section III.
IV-A Point Query
@@ IV-B Range Query
Remark 2: The range query above is indeed a rectangle range query. As for a circle range query with a center point $p=(x,y)$ and a radius $r$, a practical solution is to construct a minimal bounding rectangle (MBR) for this circle, and perform a regular rectangle range query first. Finally, a subsequent refinement step is necessary to filter out false positives.
IV-C $k$NN Query
@@
$\lceil\frac{\ln{\sqrt{(x_{u}-x_{l})^{2}+(y_{u}-y_{l})^{2}}}-\ln{\sqrt{\frac{k(x_{u}-x_{l})(y_{u}-y_{l})}{\pi N}}}}{\ln{\frac{4k}{\pi(k-1)}}}\rceil$ (3)
@@ IV-D Join Query
Remark 3: LiLIS also support a few other types of joins (e.g., $k$NN-join) out of the box, but how to leverage the learned index to design a unified framework for more types of joins is left as a future work.
V Experiment
In this section, we conduct extensive experiments to demonstrate the superiority of LiLIS over both real-world and synthetic datasets. All algorithms in this paper were implemented in Java, and all experiments were performed in a cluster with 7 CentOS machines on Tencent Cloud, in which each machine is shipped with 8 GB memory, 80 GB SSD and Intel Duo Processor of 3.0 GHz. As mentioned before, the design of LiLIS is independent with the specific distributed engine, and we present it on Apache Spark 3.0 using Standalone mode, in which one machine is chosen as the master node, and the remaining 6 machines are slave nodes. All configurations in Apache Spark, including garbage collection (GC) and shuffle operations, were maintained at their default settings for the experiments.
V-A Setup
V-A1 Datasets
The datasets used in the experiments are summarized in Table I. Both CHI [^2] https://data.cityofchicago.org/ and NYC [^3] https://www.nyc.gov/site/tlc/about/data.page contain extra information in addition to spatial locations, while SYN consists of spatial points only, which are generated by Spider [^4] https://spider.cs.ucr.edu/. As we can see, although it is possible to load CHI and SYN in a single-node machine, the performance in a distributed setting among different frameworks is a meaningful research question, and the evaluation is enough to demonstrate the effectiveness and efficiency of LiLIS. As for spatial join queries, we constructed 7000 polygons in Chicago, and perform a spatial join with CHI. For other queries, we use NYC as the default dataset.
TABLE I: Spatial Datasets
@@ V-A1 Datasets
Remark 4: While it is true that the default dataset (i.e., NYC) can be accommodated within the memory of a single contemporary machine, the primary objective of this study is to demonstrate the feasibility and efficacy of LiLIS within distributed cloud environments. In such settings, cost-effective commodity hardware with limited memory capacities (e.g., 8 GB), as utilized in our experiments, is typically employed, highlighting the system’s scalability and performance under resource-constrained conditions.
V-A2 Baselines and Implementations
For evaluation, we compare LiLIS with the state-of-the-art big spatial analytics Sedona [9]. Sedona is shipped with two indices (including R-tree and Quadtree) and two partitioners (including KD-tree and Quadtree), and we use Sedona-{I}{P} to denote its various implementation variants, in which {I} and {P} mean the index type and partitioner, respectively. For example, Sedona-RK uses R-tree as the index while adopts KD-tree as the partitioner. The four variants of Sedona are summarized in Table II. Notably, Sedona also allows the no-indexing solution, and we donate it as Sedona-N. As for the implementation using vanilla Apache Spark, we donate it as Spark. As discussed in Section III, LiLIS can adopt different partitioners, we use LiLIS-F, LiLIS-A, LiLIS-Q, LiLIS-K, and LiLIS-R to denote the implementations with fixed-grid, adaptive-grid, quadtree, KD-tree, and R-tree, respectively. In our experiments, we adopt the KD-tree as the default partitioning method for LiLIS. Consequently, LiLIS-K serves as the default implementation in our proposed framework.
TABLE II: Sedona Variants